# CKD: Contrastive Knowledge Distillation from A Sample-wise Perspective

### v1

@@ IV-A5 Performance on ImageNet-1K
To further verify the ability of CKD, we perform experiments on the ImageNet-1K dataset. Table III and Table IV report experimental results on the ImageNet-1K dataset. We can see that our method achieves superior performance with both teacher-student architectures. Moreover, the performance of CKD surpasses state-of-the-art results, providing further evidence of the effectiveness of our approach. Specifically, the proposed CKD achieves accuracy improvements of 1.58% and 4.39% over vanilla KD with the same and different teacher-student architectures.
IV-B Object Detection
@@ IV-B1 Datasets
We conduct experiments on the MS-COCO dataset [43]. MS-COCO is a popular dataset featuring 80 categories, divided into two subsets, including a training set of 118,000 images and a validation set of 5,000 images.
@@ IV-B2 Experimental Setup
We analyze ResNet-101 & ResNet-18, ResNet-101 & ResNet-50, and ResNet-50 & MobileNetV2 teacher-student pairs. Moreover, we evaluate the MS-COCO validation set with the baseline provided by Detectron2.
@@ IV-B3 Implementation Details
Experiments are performed on NVIDIA RTX 3090. We utilize Faster R-CNN with FPN as the feature extractor and implement CKD on the R-CNN head. We set $\alpha$ to 0.1 and the temperature $\tau$ to 1.0. Schedulers and task-specific loss settings are following Detectron2.
IV-B4 Performance on MS-COCO
Table V presents the object detection results on the MS-COCO dataset under three teacher-student settings. We can see that the proposed approach obtains comparable performance against logit-based approaches but inferior performance compared to feature-based approaches, particularly for ReviewKD. Object detection relies on spatial representations to locate objects. However, logits lack detailed spatial information [36, 76]. ReviewKD [7] conducts multi-level feature distillation, resulting in better performance. We combine ReviewKD with logit-based approaches to utilize spatial information. Notably, the combined approaches, including CKD+ReviewKD and DKD+ReviewKD, significantly improve object detection results compared to previous counterparts.
@@ IV-C Ablation Study
IV-C1 Effects of Temperature $\tau$
In the training phase, a high $\tau$ yields a smooth distribution across classes, making distillation more challenging. But, a low $\tau$ leads to a sharp distribution. To investigate the effect of $\tau$ on CKD, we conduct ablation studies by using ResNet32×$\times 4$ ResNet8×$\times 4$ and ResNet32×$\times 4$ ShuffleNetV1 pairs. Typically, $\tau$ is set to a low value within a range of 0.07 to 4.0, while other parameters are kept constant.
TABLE VIII: Classification accuracy (%) on the CIFAR-100 validation set. WRN-16-2 is used as a student. In the left-hand section of the table, WRN series networks serve as teachers, while in the right-hand section, different series networks are utilized as teachers.
Methods WRN Series Different Series
Teacher WRN-40-2 WRN-16-4 WRN-28-4 VGG13 WRN-16-4 ResNet50
@@ IV-C1 Effects of Temperature $\tau$
Fig. 7 presents Top-1 classification accuracies for different temperatures $\tau$. We can observe that our method performs best when temperature $\tau$ is around 1.0 and performs poorly when $\tau$ is low or high. This suggests that our method is better suited for aligning ${t}_{i}$ and ${s}_{i}$ when the value of $\tau$ is near 1.0.
From Theorem 1, we conclude that the gradient $\nabla_{\boldsymbol{s}_{*}}\mathcal{L}_{i}$ is proportional to $g_{i}$. When $\tau<1$, the difference between $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}$ and $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}$ is magnified exponentially. In the limit, we obtain $\lim_{\tau\to 0}g_{i}=0$. As a result, $g_{i}$ approaches zero, causing a rapid decline in performance. When $\tau>1$, the difference is lessened and $g_{i}$ becomes more sensitive to negative samples as $\lim_{\tau\to\infty}g_{i}=\frac{B}{B+1}$.
IV-C2 Effects of Batchsize
To analyze the effect of batch size, we utilize two teacher-student pairs including ResNet56 & ResNet20 and ResNet32×$\times 4$ ShuffleNetV2. Moreover, we set the batch size to $2^{k}$, where $k\in[4,11]\cap k\in\mathbb{N}$.
Fig. 8 presents Top-1 classification performance with varying batch sizes. We can see that our method attains good performance when the batch size is 64 for ResNet56 & ResNet20 and 32 for ResNet32×$\times 4$ ShuffleNetV2. Larger or smaller batch sizes are not optimal.
We also find that a large batch size potentially hampers the efficacy of our method as it affects intra-sample alignment through an increased $\beta$. However, a small batch size can lead to insufficient gradient magnitude. In fact, the optimal batch size for the sample-wise formulation is strongly correlated with the number of categories. It should be smaller than the number of categories to avoid neglecting intra-sample distillation and facilitate effective inter-sample distillation.
@@ IV-C3 Influence of Contrastive Formulation
Fig. 6 visualizes class-wise and sample-wise contrastive formulations. SupCon [28] is a class-wise approach that selects positive pairs within the same class. Noat that CRD [70] is a typical class-wise method with positive pairs from the same sample and negative pairs from different classes. In contrast, CKD chooses negative pairs from different samples but potentially from the same class.
Table VI presents classification performance under different formulations. We can observe that CKD outperforms SupCon and CRD because it uses sample-wise negative samples. Moreover, CRD exploits a memory bank to access a large number of negative samples and achieves a higher accuracy than SupCon.
@@ IV-C4 Influence of Triple Design
In knowledge distillation, the element of triples can be sampled from teacher or student models. Hence, there are several kinds of triples, including $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ and $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ means that we align $\mathbf{s}_{i}$ and $\boldsymbol{t}_{i}$ while distancing $\boldsymbol{s}_{i}$ from $\boldsymbol{s}_{j}$. This triple can be implemented by modifying diagonal elements of the affinity matrix with $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$. Due to fixed teacher logits, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ equals intra-sample alignment.
Table VII shows the classification performance with different triples. We can see that the utilization of triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ yields better performance than the other triples. Our method anchors on $\boldsymbol{t}_{i}$, pulling $\boldsymbol{s}_{i}$ closer while pushing $\boldsymbol{s}_{j}$ apart. This objective is clear and simple. However, as discussed in Section III-B, triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflict. For triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, this strategy mainly performs contrastive learning in the student logits space.
@@ IV-D1 Training Efficiency
To evaluate the training efficiency of our approach, we compare its accuracy and parameter count against state-of-the-art methods on the CIFAR-100 dataset. Fig. 9 depicts training time per batch and classification performance. We can observe that our method exhibits high training efficiency, striking a balance between model parameters and classification performance. Typically, feature-based methods require more parameters and longer training times than logit-based methods. ReviewKD and CRD, in particular, take twice the training time compared to our method, due to multi-level feature review and memory bank usage.
IV-D2 Effects of Teacher Capacity
Previous works [10, 74] have found that large teacher models may not be optimal for smaller student models due to disparities in parameters and capacity. Therefore, we further conduct experiments to analyze the influence of teacher networks with different capacities.
Table VIII presents classification performance on the CIFAR-100 dataset. We can see that our method achieves comparable performance across different WRN models, whereas vanilla KD exhibits variation in performance. Wide Residual Networks [73], is a ResNet variant. Therefore, our method also obtains an accuracy on par with ResNet50 as the teacher model, while vanilla KD has the second-worst performance. For vanilla KD, larger models lead to performance degeneration, for example, the accuracies of WRN-16-4 and WRN-28-4 are 76.00 and 75.14, respectively. Both the proposed method and vanilla KD perform poorly with the VGG13 architecture. Ablation studies prove that our method can partially mitigate the parameter and capacity gaps of larger models.
@@ IV-E Visualizations
GradCAM identifies image regions critical for concept prediction. To evaluate the consistency of CKD with teacher networks, we employ GradCAM to visualize class activation maps in Fig. 10. The teacher-student pair, including ResNet34 and ResNet18, are trained on the ImageNet-1K dataset. We can observe that our method effectively handles complex scenes and multiple objects. For example, in the seventh column, two birds are present. Our method accurately replicates teacher knowledge by focusing on both birds, whereas student networks and vanilla KD focus only on the bird closer to the camera.
@@ V Conclusion
In this paper, we have presented a contrastive knowledge distillation approach for image classification and object detection. To be specific, we formulate knowledge distillation as a logit alignment problem, mimicking the intra-sample numerical value and inter-sample semantic structure of teacher models. By converting the structure-preserving alignment into a sample-wise contrastive learning framework, our approach can be effectively and efficiently trained with well-designed positive and negative pairs. Experimental results demonstrate that our method achieves highly competitive performance against state-of-the-art methods on benchmark datasets. In the future, we will attempt to apply the proposed method to distill large models.

### v2

@@ IV-A5 Performance on ImageNet-1K
To further verify the effectiveness of CKD on a large-scale dataset, we perform experiments on ImageNet-1K. As listed in Table II, our method achieves superior performance with both teacher-student architectures, providing further evidence of the effectiveness on large-scale datasets. Specifically, when distilling between homogeneous networks, CKD achieves 72.24% Top-1 accuracy, representing a 1.58% improvement over vanilla KD. More remarkably, in the cross-architecture setting, CKD attains 72.97% Top-1 accuracy with a substantial 4.39% gain. These results not only validate its compatibility with diverse network structures but also establish new state-of-the-art performance on large-scale knowledge distillation.
IV-A6 Performance on Places365
To evaluate our method on large-scale scene recognition, we conduct experiments on the Places365 validation set. As shown in Table III, our method achieves the SOTA Top-1 accuracy and competitive Top-5 accuracy, demonstrating its effectiveness in complex scene understanding tasks. Notably, our method outperforms CRD in all settings, highlighting the superiority of sample-wise contrastive learning over class-wise approaches. These results validate the robustness and scalability of our method in large-scale knowledge distillation tasks.
IV-B Object Detection and Instance Segmentation
@@ IV-B1 Datasets
We evaluate our method on the MS-COCO dataset [43], which contains 118,000 training and 5,000 validation images across 80 object categories.
Figure 9: Comparison of training time per batch and accuracy on CIFAR-100. The teacher model is ResNet32$\times$4, and the student model is ResNet8$\times$4. Additional parameters are denoted after $+$ and are added based on the student’s foundation.
@@ IV-B2 Experimental Setup
We analyze ResNet-101 & ResNet-18, ResNet-101 & ResNet-50, and ResNet-50 & MobileNetV2 teacher-student pairs. Following standard evaluation protocols, we implement the detection framework using Detectron2.
@@ IV-B3 Implementation Details
All experiments are conducted on NVIDIA RTX 3090 GPUs. For object detection, we employ Faster R-CNN with FPN as the feature extractor and implement CKD on the R-CNN head. For instance segmentation, we utilize Mask R-CNN as the baseline architecture. We set $\alpha$ to 0.1 and the temperature $\tau$ to 1.0. All training protocols, including learning rate schedulers and loss functions, strictly follow Detectron2.
IV-B4 Object Detection Performance on MS-COCO
Table IV presents the object detection results on the MS-COCO dataset under three teacher-student settings. We can see that the proposed CKD achieves comparable performance against logit-based approaches but inferior performance compared to feature-based approaches, particularly for ReviewKD. This can be attributed to the fact that object detection heavily relies on spatial representations to locate objects. However, logits lack detailed spatial information [36, 78]. ReviewKD [7] conducts multi-level feature distillation, leading to superior performance. To leverage spatial information, we combine ReviewKD with logit-based approaches. Notably, the combined approaches, including CKD+ReviewKD and DKD+ReviewKD, significantly improve detection results compared to their standalone counterparts.
IV-B5 Instance Segmentation Performance on MS-COCO
Table V presents instance segmentation results on the MS-COCO dataset across different frameworks. Notably, existing knowledge distillation methods specifically designed for instance segmentation remain scarce, which limits comparisons primarily to ReviewKD and the proposed CKD+ReviewKD. CKD+ReviewKD achieves performance gains over ReviewKD, with absolute accuracy increases of +0.97, +0.91, and +1.24 points in different teacher-student pairs, respectively. The most significant improvement emerges in the RN50-MobileNetV2 setting, particularly for small objects. The synergy between logit-based distillation CKD and feature alignment ReviewKD addresses both global semantic preservation and local spatial detail transfer. This dual mechanism explains the balanced improvements across all object scales [36]. Experiments demonstrate robustness across different network gaps from RN101-RN50 to RN50-MN-V2, suggesting wide applicability in practical scenarios where compact models are essential.
@@ IV-C Ablation Study
IV-C1 Effect of Temperature $\tau$
In the training phase, a high $\tau$ yields a smooth distribution across classes, making distillation more challenging. But a low $\tau$ results in a sharp distribution. To investigate the effect of $\tau$, we conduct ablation studies by using ResNet32×$\times 4$ ResNet8×$\times 4$ and ResNet32×$\times 4$ ShuffleNetV1 pairs. Typically, $\tau$ is set to a low value within a range of 0.07 to 4.0, while other parameters are kept constant.
TABLE VIII: Classification accuracy (%) on the CIFAR-100 validation set. WRN-16-2 is used as a student. In the left-hand section, WRN series networks serve as teachers, while in the right-hand section, different series networks are utilized as teachers.
Method WRN Series Different Series
Teacher W-40-2 W-16-4 W-28-4 VGG13 W-16-4 RN50
@@ IV-C1 Effect of Temperature $\tau$
Fig. 7 illustrates the temperature sensitivity analysis for distillation. Both settings exhibit a clear performance parabola with peak Top-1 accuracy 79.0% and 76.3%, respectively, at optimal temperature $\tau=1.0$. Our method achieves optimal performance when temperature $\tau$ is approximately 1.0 and performs poorly when $\tau$ is low or high. This suggests that proper temperature scaling balances the sharpness of knowledge transfer.
From Theorem 1, we conclude that the gradient $\nabla_{\boldsymbol{s}_{*}}\mathcal{L}_{i}$ is proportional to $g_{i}$. When $\tau<1$, the difference between $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}$ and $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}$ is exponentially magnified. In the limit, we obtain $\lim_{\tau\to 0}g_{i}=0$. As a result, $g_{i}$ approaches zero, causing a vanishing gradient. When $\tau>1$, the difference is reduced and $g_{i}$ becomes more sensitive to negative samples, as $\lim_{\tau\to\infty}g_{i}=\frac{B}{B+1}$.
@@
IV-C2 Effect of Batchsize
To analyze the effect of batch size, we utilize two teacher-student pairs: ResNet56 & ResNet20 and ResNet32×$\times 4$ ShuffleNetV2. We set the batch size $B$ to $2^{k}$, where $k\in[4,11]\cap k\in\mathbb{N}$.
Fig. 8 presents top-1 accuray across various batch sizes. We can see that our method achieves maximum accuracies when the batch size is 64 for ResNet56 & ResNet20 and 32 for ResNet32×$\times 4$ ShuffleNetV2. Batch sizes outside these optimal values yield suboptimal results.
Experiments reveal that a large batch size potentially hampers the efficacy of our method by overwhelming inter-sample distillation through an increased $\beta$. However, a small batch size suffers from insufficient negative pairs due to vanishing gradients. The empirical optima should satisfy the category count constraint $B<c$. When $B>c$, distillation focuses on inter-sample relations, degrading class-specific knowledge transfer. This analysis proves that batch size significantly balances intra/inter-sample knowledge transfer.
@@ IV-C3 Influence of Contrastive Formulation
Fig. 6 visualizes differences between class-wise and sample-wise contrastive learning frameworks. In class-wise approaches such as SupCon [28], positive pairs are formed exclusively within the same class. This category also includes methods like CRD [72], which specifically constructs positive pairs from teacher-student models of the same sample while selecting negative pairs from different classes. By contrast, the sample-wise CKD adopts a different strategy in that it selects negative pairs from distinct samples, even when they potentially belong to the same class.
Table VI compares classification performance across three contrastive learning frameworks. Notably, CRD exhibits improved accuracy over SupCon by maintaining a memory bank to access negative instances. However, our CKD achieves superior performance by employing sample-wise negative pairs, which is particularly evident in heterogeneous settings where it outperforms CRD by 2.53% with the SN-v2 setting.
@@ IV-C4 Influence of Triple Design
In knowledge distillation, triplet elements can be sampled from either the teacher (t) or student (s) models. Hence, there are four kinds of triples, including $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ and $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ means that we align $\mathbf{s}_{i}$ and $\boldsymbol{t}_{i}$ while distancing $\boldsymbol{s}_{i}$ from $\boldsymbol{s}_{j}$. This triple is implemented through diagonal modification of the affinity matrix between $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$ pairs. Notably, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ reduces to intra-sample alignment since teacher logits remain fixed during distillation..
Table VII shows the classification performance for three triple settings. The teacher-anchored triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ demonstrates superior performance across all settings, achieving 1.43% higher accuracy than $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ in SN-V2 heterogeneous architecture. This advantage stems from its clear optimization objective by anchoring on fixed teacher features $\boldsymbol{t}_{i}$ to simultaneously attract student $\boldsymbol{s}_{i}$ and repel $\boldsymbol{s}_{j}$. However, as discussed in Section III-B, triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ risks gradient conflict due to competing objectives between moving $\boldsymbol{s}_{i}$ toward $\boldsymbol{t}_{i}$ and away from $\boldsymbol{t}_{j}$. For triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, this strategy mainly performs contrastive learning in the student logits space, yet insufficiently leverages the prior knowledge from the teacher model.
@@ IV-D1 Training Efficiency
To evaluate the efficiency-accuracy trade-off, we compare our method against knowledge distillation baselines on CIFAR-100. As shown in Fig. 9, our method exhibits high training efficiency, striking a good balance between model parameters and classification performance. Feature-based methods require more parameters and longer training times than logit-based methods. ReviewKD and CRD, in particular, take twice the training time over our method due to multi-level feature review and memory bank. Our method simultaneously improves accuracy and reduces training time.
IV-D2 Effect of Teacher Capacity
Existing studies [10, 76] have demonstrated that large teacher models may be suboptimal for distilling knowledge into smaller student models due to parameter and capacity gaps between the two architectures. To investigate this phenomenon, we conduct experiments to examine how teacher networks with varying model capacities affect knowledge transfer efficiency.
Table VIII compares knowledge distillation performance on the CIFAR-100 dataset. We can see that our method achieves comparable performance with increasing WRN teacher capacity, whereas vanilla KD exhibits variation in performance. For vanilla KD, larger models lead to performance degradation, achieving the accuracies of WRN-16-4 and WRN-28-4 at 76.00 and 75.14, respectively. In contrast, our CKD achieves progressive accuracy improvements as teacher capacity increases from WRN-40-2 to WRN-28-4, demonstrating effective utilization of enhanced teacher representations. Wide Residual Networks [75] is a ResNet variant. Our method also obtains an accuracy on par with ResNet50 as the teacher model, while vanilla KD has the second-worst performance. Both methods show limited effectiveness with VGG13 architecture, suggesting architectural compatibility influences distillation efficacy. These results prove that our method can partially mitigate the parameter and capacity gaps of larger models.
@@ IV-E Visualizations
Fig. 10 demonstrates the semantic alignment between CKD and teacher networks through GradCAM visualizations, where the ResNet34-ResNet18 pair was trained on ImageNet-1K. We can observe that our method effectively handles complex scenes and multiple objects. In multi-object cases, CKD captures both discriminative features of birds circled in red, while the student baseline and vanilla KD focus solely on the foreground bird. Across diverse scenes like natural landscapes and indoor settings, CKD consistently preserves the spatial attention patterns, particularly in fine-grained regions like butterfly wing patterns. In the final column, CKD aligns with the teacher model in localizing the target dog, whereas the student network and vanilla KD erroneously emphasize the distractor dog. These visual evidence confirm that CKD maintains teacher-aligned sample-specific semantics rather than generic category features.
@@ V Conclusion
In this paper, we have presented a contrastive knowledge distillation approach for image classification and object detection. To be specific, we formulate knowledge distillation as a logit alignment problem, mimicking the intra-sample numerical value and inter-sample semantic structure of teacher models. By converting the structure-preserving alignment into a sample-wise contrastive learning framework, our approach can be effectively and efficiently trained with well-designed positive and negative pairs. Experiments demonstrate that our method achieves highly competitive performance against state-of-the-art methods on benchmarks. In the future, we will explore the application of this framework to large multimodal models.